# Rewriting Pre-Training Data Boosts LLM Performance in Math and Code

### v3

@@ A.4.2 Math ablation data mixture
Appendix B Code linting filtering
A threshold of 7.0 balances code quality and dataset size, while the comment penalty reduces overly verbose or non-functional scripts. As described in Section 3.2.2, the code for performing linter filtering is publicly available at https://github.com/rioyokotalab/swallow-code-math. The relevant code is also provided below.
When conducting linting with pylint, warnings and errors dependent on the linting environment, such as import errors, are excluded using the --disable option. Additionally, some files with Python extensions primarily contain comments with textual content and have minimal script functionality. To address this, as discussed in Section 3.2.2, we introduced a mechanism that imposes a heuristic penalty based on the proportion of comments to filter out such files.
@@ Appendix B Code linting filtering
Appendix C Code LLM-based scoring
As described in Section 3.2.3, this section presents the prompt used for LLM-based scoring. The prompt was provided to Llama-3.3-70B-Instruct to evaluate code quality. The scoring criteria were developed with reference to the Google Python Style Guide  [^13] https://google.github.io/styleguide/pyguide.html. The actual code pipeline is available at https://github.com/rioyokotalab/swallow-code-math. The dataset created through this process is available at https://huggingface.co/datasets/tokyotech-llm/swallow-code, and the checkpoints of the model trained on this dataset are available at https://huggingface.co/collections/tokyotech-llm/swallowcode-6811c84ff647568547d4e443.
@@ 10. Is the code formatted in a way that enhances readability?
Figure 8: Distribution of quality scores assigned by Llama-3.3-70B-Instruct to the Python code corpus. The dashed line at 6 denotes the filtering threshold. Notably, only even scores appear, an emergent property of how Llama-3.3-70B-Instruct interpreted the scoring prompts.
Figure 8 illustrates the distribution of the quality scores assigned.
Appendix D Code LLM rewriting
@@
D.1 Style-Guided Code Rewriting (SGCR)
The prompt used for SGCR is provided below. The dataset created through this rewriting process is available at https://huggingface.co/datasets/tokyotech-llm/swallow-code, and the checkpoints of the model trained on this dataset are available at https://huggingface.co/collections/tokyotech-llm/swallowcode-6811c84ff647568547d4e443.
@@ 10. Is the code formatted in a way that enhances readability?
D.2 Self-Contained Optimization Rewriting (SCOR).
The prompt used for SCOR is provided below. The dataset created through this rewriting process is available at https://huggingface.co/datasets/tokyotech-llm/swallow-code, and the checkpoints of the model trained on this dataset are available at https://huggingface.co/collections/tokyotech-llm/swallowcode-6811c84ff647568547d4e443.
@@
Appendix E Math LLM rewriting
@@
Appendix F Computational cost
This section quantifies the computational cost, measured in GPU hours (H100), required for the LLM scoring, LLM rewriting, and ablation experiments conducted in this study. These measurements are based on empirical data. By analyzing the computational cost, we aim to elucidate the relationship between the resource demands of the data cleaning pipeline and the resulting improvements in downstream task performance for each method. Furthermore, by providing an estimate of the computational resources needed to reproduce our results, we seek to facilitate future research.
F.1 Computational cost of LLM scoring
@@
The data synthesis process utilized vLLM 0.7.2 and PyTorch 2.5.1. The dataset was synthesized with a global batch size of 2048 and tensor parallelism of 4. Data generation was performed using four H100 (94 GB) GPUs per job. With an input processing speed of approximately 2000 tokens/s and an output generation speed of approximately 3000 tokens/s, and given an average input dataset length of 836 tokens, an average output dataset length of 1271 tokens, and a total sample count of 20,826,548, we estimate that the dataset creation for the experiments described in Section 3.2.3 consumed a total of 19,477 H100 GPU hours. This estimate excludes vLLM initialization and safetensor loading.
F.2 Computational cost of LLM rewriting
@@
The synthesis process for LLM rewriting also employed vLLM 0.7.2 and PyTorch 2.5.1, with a global batch size of 2048 and tensor parallelism of 4. Data generation was conducted using 4 H100 (94GB) GPUs per job. The input processing speed was approximately 2000 tokens/s, and the output generation speed was approximately 3000 tokens/s. Given an average input dataset length of 836 tokens, an average output dataset length of 1819 tokens, and a total sample count of 20,826,548, we estimate that the dataset creation for the experiments described in Section 3.3.1 consumed a total of 23,703 H100 GPU hours. This estimate excludes vLLM initialization and safetensor loading.
F.3 Computational cost of continual pre-training data ablation experiments
@@
The ablation experiments described in Sections 3.1 and 4.1 were conducted using 64 H100 (94GB) GPUs for 24.7 hours per experiment. Thus, each experiment consumed 1,580 H100 GPU hours. With a total of 15 experiments (13 code ablation experiments and two math ablation experiments), the overall computational cost amounted to 23,700 H100 GPU hours. These experiments achieved a training throughput of 530 TFLOP/s/GPU and 590,000 tokens/s, as derived from the FLOP/s formula presented in the Megatron-LM paper [37]. This corresponds to approximately 53.5% of the peak BF16 Tensor Core performance of the H100.
Appendix G MBPP
@@ Appendix H Evaluations
H.1 Code ablation experiments results
As described in Section 3.1, we evaluated models continually pre-trained from Llama-3.1-8B on ten English downstream tasks. In the following, we report the evaluation results for 13 code ablation experiments, with their relationships illustrated in Figure 10. Experiments exp1, exp8, exp9, and exp13 serve as baselines, utilizing only Python data extracted from existing open code corpora. The remaining experiments are conducted to construct the SwallowCode dataset. We evaluated performance using the following ten benchmarks: OpenBookQA [22], TriviaQA [23], HellaSwag [24], SQuAD 2.0 [25], XWinograd [26], MMLU [27], GSM8K [28], BBH [29], HumanEval [30], and HumanEval+ [31].
@@ H.1 Code ablation experiments results
H.2 Math ablation experiments results
As described in Section 4.1, we evaluated models continually pre-trained from Llama-3.1-8B on ten English downstream tasks. Below, we present the evaluation results for two math ablation experiments. Specifically, we adopted the following ten evaluation benchmarks: OpenBookQA [22], TriviaQA [23], HellaSwag [24], SQuAD 2.0 [25], XWinograd [26], MMLU [27], GSM8K [28], BBH [29], HumanEval [30], and MATH [27].
@@ H.2 Math ablation experiments results
Experiment 2: finemath-4+ rewriten(Llama-3.3-70B-Instruct)

### v4

@@ A.4.2 Math ablation data mixture
  • Japanese Wikipedia: 0.84 billion tokens
  • Japanese Swallow Corpus v2: 33.0 billion tokens
  • Laboro-ParaCorpus: 0.22 billion tokens
  • English Wikipedia: 1.1 billion tokens
  • English Cosmopedia: 3.3 billion tokens
  • English DCLM: 2.2 billion tokens
  • SwallowCode (Syntax, Pylint Filtered): 6.5 billion tokens
  • Target Math Data: 2.4 billion tokens
A.5 Impact of Language Mixture on Ablation Experiments
Our ablation experiments utilize a data mixture with a high proportion of Japanese text, which is controlled for fair comparisons but represents an unconventional setup for English-centric benchmarks like HumanEval. To investigate potential subtle effects on learning from the code and math portions, we performed additional experiments using an alternative mixture: 70% English text (sourced from DataComp-LM and Cosmopedia), 20% code (SwallowCode), and 10% math (SwallowMath). The experimental setup otherwise follows that described in Section 3.1, with continual pre-training of Llama-3.1-8B for 50B tokens. The results are summarized in Table 6.
Table 6: Performance on Key Benchmarks with English-Heavy Mixture
Benchmark GSM8K MATH HumanEval HumanEval+
Accuracy 0.700 0.354 0.583 0.540
These scores are higher on English-centric math and code benchmarks compared to our original Japanese-heavy mixture, suggesting that a high Japanese proportion may diminish performance on such tasks within a limited 50B-token budget, possibly due to reduced exposure to English-aligned patterns. However, since the original experiments in Sections 3.1 and 4.1 are controlled, the improvements of SwallowCode and SwallowMath remain robust and unaffected by the composition of the mixture.
Appendix B Generalizability beyond Llama-3: Qwen2-7B
To assess generalizability beyond the Llama-3 family, we conduct 20B-token continual pre-training starting from Qwen2-7B. We use the final SwallowCode (post-SCOR) as the code component and adopt an alternative mixture of 70% English text (DataComp-LM and Cosmopedia), 20% code, and 10% math (SwallowMath). Unless otherwise noted, all hyperparameters and evaluation benchmarks follow the protocol in Section 3.1 and Appendix J, except that we use LR=1.0E-5, a global batch size of 1,024, and a sequence length of 4,096. As a code baseline, we replace SwallowCode with Stack-Edu under the same token budget and mixture ratios. Compared to Stack-Edu, SwallowCode yields +10.3 pass@1 on HumanEval and +10.3 pass@1 on HumanEval+, reaching 49.6 and 44.6 respectively (Table 7). These results indicate that the benefits of our transform-and-retain pipeline are not specific to Llama-3–based models.
Table 7: Qwen2-7B, 20B-token continual pre-training. Scores are pass@1.
Corpus HumanEval HumanEval+
Stack-Edu 39.3 34.3
SwallowCode (ours) 49.6  ( +10.3) 44.6  ( +10.3)
Appendix C Code linting filtering
A threshold of 7.0 balances code quality and dataset size, while the comment penalty reduces overly verbose or non-functional scripts. As described in Section 3.2.2, the code for performing linter filtering is publicly available. The relevant code is also provided below.
When pylint is applied, warnings and errors dependent on the linting environment, such as import errors, are excluded using the --disable option. Additionally, some files with Python extensions contain primarily comments with textual content and have minimal script functionality. To address this, as discussed in Section 3.2.2, we introduced a mechanism that imposes a heuristic penalty based on the proportion of comments to filter out such files.
@@ Appendix C Code linting filtering
Appendix D Code LLM-based scoring
As described in Section 3.2.3, this section presents the prompt used for LLM-based scoring. The prompt was provided to Llama-3.3-70B-Instruct to evaluate code quality. The scoring criteria were developed with reference to the Google Python Style Guide [^11] https://google.github.io/styleguide/pyguide.html.
@@ 10. Is the code formatted in a way that enhances readability?
Appendix E Code LLM rewriting
@@
E.1 Token Retention Analysis
The rewriting processes in SGCR and SCOR inherently modify the token lengths of data samples. To provide insights into these transformations and their implications for training efficiency, we computed the average input and output token lengths across the dataset, as summarized in Table 8.
Table 8: Average Input and Output Token Lengths for Rewriting Methods
Rewriting Method Input Tokens Output Tokens
SGCR 836 548
SCOR 548 835
For SGCR, output lengths are shorter than input lengths on average due to the pipeline discarding samples where the original code combined with the prompt and generated text exceeds the model’s maximum context length, or where generated code cannot be reliably extracted (e.g., absence of a code block like ‘‘‘python) Our analysis shows that SGCR generally condenses the data. In contrast, SCOR expands the data.
E.2 Syntax Error Analysis
While LLM rewriting can potentially introduce syntax errors or other implementation issues, constructing a dataset at the billion-token scale renders it infeasible to detect and remove all such anomalies comprehensively. To provide insights into the extent of code error accumulation during the LLM rewriting process, we measured the syntax error rates after each rewriting stage on a random subset of 100,000 samples, finding rates of 0.73% post-SGCR and 0.46% post-SCOR. These results indicate that errors do not accumulate across the two-stage rewriting process, alleviating concerns about progressive degradation. Furthermore, incorporating syntax error filtering at each rewriting step holds promise for enhancing dataset quality in future works.
E.3 Rationale for the Two-Stage Rewriting Pipeline
A natural question is whether the SGCR and SCOR stages could be merged into a single rewriting pass. In preliminary experiments, we combined all instructions from both prompts into a unified prompt containing 19 distinct directives. Manual inspection of the outputs revealed that Llama-3.3-70B-Instruct frequently exhibited instruction drift: the model complied with only a subset of the directives while neglecting others. For example, when style-related and self-containment instructions were presented together, the model tended to prioritize structural changes (e.g., making the code self-contained) at the expense of stylistic refinements (e.g., docstrings, type annotations), or vice versa.
To ensure robust adherence to both stylistic and semantic requirements, we adopted the decoupled two-stage pipeline described in Section 3.3. This decomposition allows each stage to target a focused set of objectives, reducing the cognitive load on the rewriting model and yielding more consistent outputs.
E.4 Style-Guided Code Rewriting (SGCR)
The prompt used for SGCR is provided below.
@@ 10. Is the code formatted in a way that enhances readability?
E.5 Self-Contained Optimization Rewriting (SCOR).
The prompt used for SCOR is provided below.
@@
Appendix F Math LLM rewriting
@@
Appendix G Computational cost
This section quantifies the computational resources, measured in H100 GPU hours, required for the LLM scoring, LLM rewriting, and ablation experiments in this study. These estimates are derived from empirical measurements. By detailing these costs, we aim to illustrate the trade-offs between the resource requirements of our data refinement pipeline and the resulting enhancements in downstream task performance. Additionally, we provide these figures to support the reproducibility of our results and inform future research efforts. Importantly, both the syntax-error filtering and the pylint-based filtering stages are CPU-only; although the dataset scale implies nontrivial processing time, their cost is negligible compared to the GPU-bound rewriting stage, which dominates the overall compute budget.
G.1 Computational cost of LLM scoring
@@
The LLM scoring process employed vLLM 0.7.2 and PyTorch 2.5.1, with a global batch size of 2048 and tensor parallelism of 4. Data generation utilized four H100 (94 GB) GPUs per job. At an input processing speed of approximately 2000 tokens/s and an output generation speed of approximately 3000 tokens/s, with an average input length of 836 tokens, an average output length of 1271 tokens, and a total of 20,826,548 samples, we estimate that the dataset creation for the experiments in Section 3.2.3 required 19,477 H100 GPU hours. This figure excludes vLLM initialization and safetensor loading times.
G.2 Computational cost of LLM rewriting
@@
Similarly, the LLM rewriting synthesis used vLLM 0.7.2 and PyTorch 2.5.1, configured with a global batch size of 2048 and tensor parallelism of 4. Generation was performed on four H100 (94 GB) GPUs per job, achieving an input processing speed of about 2000 tokens/s and an output generation speed of about 3000 tokens/s. With an average input length of 836 tokens, an average output length of 1819 tokens, and 20,826,548 samples in total, the dataset creation for the experiments in Section 3.3.1 is estimated to have consumed 23,703 H100 GPU hours, excluding vLLM initialization and safetensor loading.
G.3 Computational cost of continual pre-training data ablation experiments
@@
The ablation experiments outlined in Sections 3.1 and 4.1 were executed using 64 H100 (94 GB) GPUs for 24.7 hours per run, resulting in 1,580 H100 GPU hours per experiment. Across 15 experiments (13 for code ablations and 2 for math ablations), the total computational expenditure was 23,700 H100 GPU hours. These runs achieved a training throughput of 530 TFLOP/s/GPU and 590,000 tokens/s, calculated using the FLOP/s formula from the Megatron-LM paper (Narayanan et al., 2021). This represents approximately 53.5% of the H100’s peak BF16 Tensor Core performance.
G.4 Cost-Benefit Analysis
While rewriting at the full 16.1B-token scale for SwallowCode demands substantial GPU hours, our approach demonstrates clear efficiency gains in downstream training. As illustrated in Figure 1, continual pre-training on unrefined or noisy datasets like The-Stack-v2 or Stack-Edu results in only marginal improvements on benchmarks such as HumanEval and HumanEval+. In contrast, our rewritten data achieves significant performance boosts with equivalent or fewer tokens, highlighting enhanced data efficiency. Although we did not extend baseline experiments beyond 50B tokens, extrapolating the trends in Figure 1 indicates that unprocessed datasets would struggle to match our rewriting method’s performance, even at scales exceeding 100B tokens. To enhance practicality for teams with limited computational resources, our pipeline is modular and scalable, allowing selective application to data subsets or direct use of our publicly released datasets, thereby bypassing the rewriting overhead. This one-time investment in data refinement produces high-quality, reusable corpora that can be shared across the community, amortizing costs over multiple applications and studies. Unlike transient expenditures on model training, resources allocated to data curation yield enduring assets that advance broader research. As evidenced in works like Qwen3 (Yang et al., 2025), which invested heavily in synthesizing trillions of high-quality text tokens, such commitments are standard in LLM development and deliver substantial long-term value.
Appendix H Contamination Analysis
In this section, we detail our efforts to mitigate and verify the absence of data contamination in the SwallowCode and SwallowMath corpora. Contamination, whether through direct test-set leakage or indirect self-contamination introduced by the rewriting LLM, can inflate performance metrics and undermine the validity of experimental results. We address both concerns systematically to ensure the integrity of our findings.
H.1 Decontamination Against Evaluation Benchmarks
To prevent test-set leakage, we performed rigorous decontamination checks on both corpora against their respective downstream benchmarks. For SwallowCode, we streamed the entire 16.1B token corpus and scanned for overlaps with HumanEval and HumanEval+ prompts. This involved checking for exact matches and computing Jaccard similarity scores, with a threshold of $\geq$ 0.8 considered indicative of high similarity. No such instances were detected, confirming that SwallowCode is free of direct contamination from these benchmarks. Similarly, for SwallowMath, we conducted an analogous procedure against the prompts and solutions from GSM8K and MATH. Using exact match detection and Jaccard similarity ($\geq$ 0.8 threshold), we verified that there are no contamination overlaps in the corpus.
H.2 Addressing Self-Contamination from the Rewriting LLM
A subtler risk arises from ”self-contamination,” where the rewriting of LLM, Llama-3.3-70B-Instruct, might inadvertently incorporate patterns or knowledge from benchmark problems into the rewritten data. This model has a knowledge cutoff date of December 2023, meaning it lacks exposure to post-cutoff data or benchmarks. To evaluate this, we assessed the performance on GSM-Plus, a mathematics benchmark released after the cutoff date, which the rewriting LLM could not have encountered during its training. Continual pre-training of Llama-3.1-8B with Finemath-4+ yields 35.75 points on GSM-Plus (Li et al., 2024), while using SwallowMath achieves 46.52 points. This substantial improvement demonstrates that the gains from SwallowMath are attributable to enhanced data quality rather than embedded contamination from the rewriting process. These analyses collectively affirm that our performance improvements stem from genuine advancements in data refinement, rather than artifacts of contamination.
Appendix I MBPP
@@
J.1 Code ablation experiments results
As described in Section 3.1, we evaluated models continually pre-trained from Llama-3.1-8B on ten English downstream tasks. In the following, we report the evaluation results for 13 code ablation experiments, with their relationships illustrated in Figure 9. Experiments exp1, exp8, exp9, and exp13 serve as baselines, utilizing only Python data extracted from existing open code corpora. The remaining experiments are conducted to construct the SwallowCode dataset. We evaluated performance using the following ten benchmarks: OpenBookQA (Mihaylov et al., 2018), TriviaQA (Joshi et al., 2017), HellaSwag (Zellers et al., 2019), SQuAD 2.0 (Rajpurkar et al., 2018), XWinograd (Tikhonov & Ryabinin, 2021), MMLU (Hendrycks et al., 2021a), GSM8K (Cobbe et al., 2021), BBH (Suzgun et al., 2022), HumanEval (Chen et al., 2021), and HumanEval+ (Liu et al., 2023).
@@
J.2 Math ablation experiments results
As described in Section 4.1, we evaluated models continually pre-trained from Llama-3.1-8B on ten English downstream tasks. Below, we present the evaluation results for two math ablation experiments. Specifically, we adopted the following ten evaluation benchmarks: OpenBookQA (Mihaylov et al., 2018), TriviaQA (Joshi et al., 2017), HellaSwag (Zellers et al., 2019), SQuAD 2.0 (Rajpurkar et al., 2018), XWinograd (Tikhonov & Ryabinin, 2021), MMLU (Hendrycks et al., 2021a), GSM8K (Cobbe et al., 2021), BBH (Suzgun et al., 2022), HumanEval (Chen et al., 2021), and MATH (Hendrycks et al., 2021a).
@@
Experiment 2: finemath-4+ rewritten(Llama-3.3-70B-Instruct)
@@
Appendix K The Use of Large Language Models
In accordance with the ICLR 2026 policy on responsible use of language models, we disclose that LLMs were used only for nonsubstantive copy editing. Specifically, LLMs were employed to (i) suggest alternative word choices, (ii) correct minor grammatical issues, and (iii) improve the clarity and fluency of sentences originally written by the authors.
No LLM was used to generate scientific content, conduct data analysis, design or execute experiments, create figures or tables, select citations, or draw conclusions. All technical ideas, methods, experiments, results, and interpretations were conceived, written, and verified by the authors.
The use of LLMs does not affect the reproducibility or integrity of the work.
[…]